OpenStudy (anonymous):

derivative Find the derivative of f(x)= -9/x at x=-8 Ive gotten this far: f(x)=-9/x f(x+h)= -9/(x+h) then plugging that into the derivative formula thing: (-9/(x+h)) - (-9/x) all over h this is where i get stuck, what do i do next

OpenStudy (anonymous):

you have a couple of choices one is to know that the derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) so the derivative of \(\frac{-9}{x}\) is \(\frac{9}{x^2}\)

OpenStudy (anonymous):

how does that apply to the formula thingie?

OpenStudy (anonymous):

the other is to compute \[\lim_{x\to -8}\frac{f(x)-f(8)}{x+8}\]

OpenStudy (anonymous):

could i use direct substitution to compute that?

OpenStudy (anonymous):

but we can do it your way if you like

OpenStudy (anonymous):

yes after some algebra

OpenStudy (anonymous):

but lets do you your way

OpenStudy (anonymous):

ok thanks man

OpenStudy (anonymous):

\[\frac{-9}{x+h}+\frac{9}{x}\] add them up

OpenStudy (anonymous):

would it become \[\frac{ -9+9(x+h) }{ x^2+h }\]

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

leave the denominator in factored form as \(x(x+h)\) which is not \(x^2+h\) it is \(x^2+xh\) but don't multiply out, just leave it

OpenStudy (anonymous):

also the numerator is wrong , it should be \[-9x+9(x+h)\]

OpenStudy (anonymous):

yea the adding part is where i get confused

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\] is how you add fractions the fact that in most stupid algebra problems you can find some other denominator is misleading

OpenStudy (anonymous):

\[\frac{ 9h }{ x(x+h) }\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now divide by the \(h\) that we have been ignoring and get \[\frac{9}{x(x+h)}\]

OpenStudy (anonymous):

what would we do next?

OpenStudy (anonymous):

take the limit as \(h\) goes to zero, which really means erase the \(h\) then replace \(x\) by \(-8\) or any other number of your choosing

OpenStudy (anonymous):

so we just plug in -8 for x and 0 for h?

OpenStudy (anonymous):

first get rid of the \(h\) and you get \[f'(x)=\frac{9}{x^2}\] as promised way above then you can compute \[f'(-8)=\frac{9}{(-8)^2}\] or \[f'(whatever)\] because now you have an explicit expression for \(f'\)

OpenStudy (anonymous):

so we end up with 9/64 as the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks SO much!!

OpenStudy (anonymous):

if this seems hard or annoying, don't worry in a week you will be able to do this in your head or maybe two weeks yw